Consider a pizza restaurant where ovens are a fixed input and workers are variable inputs. Assume labor is the only variable cost for the business. The pizza restaurant has a fixed cost of $100 per day and pays each worker $150 per day.

Fill in the blanks to complete the Marginal Physical Product of Labor column for each worker and the Marginal Cost column at each level of labor. (Hint: Marginal cost is the change in total cost divided by the change in the quantity of output. You can calculate it here by dividing the increase in total cost from hiring one more worker by the marginal physical product from hiring one more worker.)

Labor

Quantity of Output

Marginal Physical Product of Labor

Total Cost

Marginal Cost

(Workers)

(Pizzas per day)

(Pizzas per day)

(Dollars per day)

(Dollars per pizza)

0 0 100
1 12 250
2 27 400
3 47 550
4 77 700
5 102 850
6 117 1,000

True or False: As marginal physical product rises, marginal cost falls.

True

False
Answer #1

Marginal physical product of labor gives the change in output when 1 additional worker is added. SO, here, MPl for 2nd worker = 12 - 0 = 12. Similarly, we fill up the table.

Marginal cost gives the change in cost when there is a change is quantity produced. SO, here, MC for Q = 12 is = (250-100)/ (12 - 0) = 150/12 = 12.5. Similarly, we fill up the table.

Labor Quantity of Output Marginal Physical Product of Labor Total Cost Marginal Cost
0 0 100
1 12 12 250 12.5
2 27 15 400 10
3 47 20 550 7.5
4 77 30 700 5
5 102 25 850 6
6 117 15 1000 10

It is true that as marginal physical product rises, marginal cost falls